the length of a rectangle is twice its width. if the perimeter of the rectangle is 54 m, find its area
Answer by lwsshak3(11628)

the length of a rectangle is twice its width. if the perimeter of the rectangle is 54 m, find its area
***
let x=width
2x=length
2(length+width)=perimeter
2(2x+x)=54
6x=54
x=9
2x=18
area=length*width=9*18=162 sq m